OpenStudy (anonymous):

Evaluate the expression and choose the correct answer in scientific notation. (2 × 10^2) × (4 × 10^2) A. 8 × 10^4 B. 7 × 10^3 C. 8 × 10^2

Vocaloid (vocaloid):

hint: (10^2)*(10^2) = ?

OpenStudy (anonymous):

And 7 × 10^2 and 200

Vocaloid (vocaloid):

???

OpenStudy (calculusxy):

\[(2 \times 10^2) \times (4 \times 10^2)\] multiply the coefficients first \[(2 \times 4)(10^2 \times 10^2)\] and now multiply each of them \[8 \times 10^4\]

OpenStudy (anonymous):

Okay thank you

OpenStudy (calculusxy):

yw :)

OpenStudy (vikstar2.0):

yeah i thinks its either a or c